OpenStudy (anonymous):

y(greater than or equal) 2x=4 and identify a point that satisfies this inequality but does not satisfy y>2x+4

OpenStudy (compassionate):

y => 2x = 4 does not make sense, and is not an expression nor equation.

OpenStudy (ranga):

Should it be 2x + 4 instead of 2x = 4?

OpenStudy (the_fizicx99):

y ≥ 2x = 4 doesn't make sense. I think you mean to have ±

OpenStudy (compassionate):

The world may never know.

OpenStudy (the_fizicx99):

^

OpenStudy (ranga):

@msumner Is it supposed to be: \[\Large y \ge 2x + 4\]

OpenStudy (anonymous):

Im sorry i meant +4

OpenStudy (ranga):

Pick any point on the line y = 2x + 4. It will satisfy the inequality y >= 2x + 4 but not y > 2x + 4

OpenStudy (the_fizicx99):

Ranga i think you have a typo there :p

OpenStudy (anonymous):

(0,0) is what they told me to pick most of the time

OpenStudy (the_fizicx99):

That's to check the origin, weather or not to shade it

OpenStudy (ranga):

What is the typo you were referring to?

OpenStudy (the_fizicx99):

y = 2x+ 4

OpenStudy (ranga):

Yes. Any point chosen on the line y = 2x + 4 will satisfy the the inequality y >= 2x + 4 but not y > 2x + 4. For example pick x = 5, then y = (2)(5) + 4 = 14 The point (5,14) satisfies y >= 2x + 4 but not y > 2x + 4

OpenStudy (anonymous):

would the points (1,6) fit

OpenStudy (ranga):

Yes. (1,6) is another point that satisfies y >= 2x + 4 but not y > 2x + 4.

OpenStudy (ranga):

To graph the inequality y >= 2x + 4 you will first draw the line y = 2x + 4. The line will be a continuous line because all points on the line will satisfy the inequality too because of the equal to in >=. You will shade the region above the line. To graph the inequality y > 2x + 4 you will first draw the line y = 2x + 4. The line will be a dashed line because all points on the line are not included and you will shade the region above the line. The difference between these two solution sets is the former includes all points on the line whereas the latter excludes it. Therefore, all points on the line y = 2x + 4 will satisfy y >= 2x + 4 but not y > 2x + 4